Skip to main contentSkip to solution

Ten cards are numbered from 1 to 10. A and B are playing a game by drawing a card, alternately, with replacement, till one gets a prime number then the probability of B's winning if A started the game is

Solution

✅ Correct Option: 4

Primes among 1 to 10 are 2, 3, 5, 7, so p=25p=\frac25 and q=35q=\frac35. B wins on the 2nd, 4th, ... draw, giving qp+q3p+⋯=qp1−q2=(3/5)(2/5)1−9/25=6/2516/25=38qp+q^3p+\dots=\frac{qp}{1-q^2}=\frac{(3/5)(2/5)}{1-9/25}=\frac{6/25}{16/25}=\frac38.

Keyboard Shortcuts

  • Left arrow: Previous question
  • Right arrow: Next question
  • S key: Jump to solution
  • Q key: Jump to question